# In-Situ Melt Pool Characterization via Thermal Imaging for Defect Detection in Directed Energy Deposition Using Vision Transformers

### v1

@@ 1 Introduction
In recent years, significant efforts have been made to develop in-situ monitoring and characterization methods for melt pool dynamics in laser additive manufacturing (LAM), especially for identifying internal defects within directed energy deposition (DED). [7] introduced a real-time porosity detection framework in DED, leveraging Self-Organizing Map (SOM) clustering on thermal melt pool images, achieving a 96% prediction accuracy. Similarly, [8] proposed a data fusion approach combining PyroNet (CNN) and IRNet (LRCN) models to predict porosity in laser-based AM with high accuracy through decision-level fusion of pyrometer and infrared image data. [9] achieved a porosity classification accuracy of 98.44% by combining thermal history and morphological features of melt pools with a K-Nearest Neighbor (KNN) model. [10] applied a variational autoencoder (VAE) combined with Gaussian mixture modeling (GMM) and K-means clustering on high-resolution thermal images to effectively detect melt pool anomalies in Directed Energy Deposition (DED) with a reported accuracy of 94.52%. In laser wire DED, [11] developed a CNN-based approach using variants of YOLO, for real-time segmentation and statistical analysis of melt pools, enhancing monitoring accuracy. Meanwhile, [12] utilized a high-dynamic range camera with a KNN classifier to classify stability zones for defect prevention, achieving a 13% error rate. Multi-sensor fusion and data-driven models have also shown promise. For instance, [13] introduced a data fusion approach integrating multiple sensors to monitor melt pool characteristics using multiple ML models such as Support Vector Machine (SVM), K-Nearest Neighbor(Knn), Random Foresta dn Multi-perceptron layers (MLP) classifiers, achieving a high true positive rate (90%) and a low false positive rate (1%) for porosity detection. [14] presented an unsupervised online learning method for real-time defect detection in laser metal deposition (LMD) using high-speed pyrometer melt pool imaging, achieving 76% accuracy with K-means and 97% with SOM for classifying images as “healthy” or “anomalous.”In another approach, [15] developed a system for defect diagnosis in DED using a degree of irregularity (DOI) feature index, reaching diagnostic accuracies of 76.65% for normal and 96.92% for balling states, thus surpassing traditional STFT (Short Time Fourier Transform) based 2D CNN models.
More recent developments continue to push the boundaries of in-situ monitoring accuracy. [16] introduced FixConvNeXt, achieving 99.1% accuracy in real-time melt pool state identification for L-DED, along with reduced computational load, offering an efficient solution for online defect detection and control. In a pioneering acoustic-based approach, [17] achieved 89% overall accuracy, 93% for keyhole pore detection, and a 98% AUC-ROC score using CNNs for DED, demonstrating the strong potential of acoustic signals for real-time quality monitoring. Recently, [18] presented a layer-wise anomaly detection method by Gaussian Support Vector Machine classifier (SVM) classifier for DED, achieving up to 96.38% accuracy and a 95.34% F-score for thin-wall specimens, and 85.07% accuracy with a 91.56% F-score for cylindrical specimens, outperforming benchmark methods in both efficiency and precision. [19] introduced a DenseNet-39 model utilizing high-speed camera melt pool images for monitoring melt pool states in Laser-Based Directed Energy Deposition (L-DED). The model achieves 99.3% accuracy in classifying melt pool states as either stable or unstable, effectively reducing the computational burden and processing time, and highlighting its suitability for efficient in-situ process monitoring. [20] used high-speed camera melt pool video data in laser hot wire DED to detect flaws like wire dripping, arcing, oscillation, and stubbing via a ConvLSTM autoencoder. The self-supervised model accurately identifies anomalies by reconstructing, enabling effective real-time monitoring in DED. Table 1 provides a summary of the different learning methodologies and models employed for defect detection tasks through in-situ melt pool data within the DED process.
@@ 2 Experimentation & Data collection
For data collection, we conduct the experiments using a customized powder-based DED system (AMBIT Core DED, Hybrid Manufacturing Technology, TX, USA). We mount a thermal imager Pyrometer camera (ThermaViz® TV200, Stratonics) inside the printing chamber for in-situ monitoring and melt pool data capture during printing. Since our DED setup is hybrid, including both additive and CNC machines, we do an off-axis installation of the camera at a viewing angle of 57 (as shown in Figure 2). The focal length is fixed at 42 mm with a working distance of $245\pm 30$ mm. A Neutral density filter is also calibrated and used to prevent pixel saturation during the in-situ recording by the pyrometer.
In this study, we print two single-track thin-wall Inconel 718 samples, each measuring 40 mm in length and consisting of 10 layers in Figure 2, with the scanning direction reversed on every layer to create a linear reciprocating scan pattern. The variations in printing strategies and process parameters are summarized in Table 2 below. Note that, we have fixed the exposure time of the pyrometer at 0.3 ms optimized by experimentation for both sample printing. The camera records the distribution of thermal gradient of the traveling melt pool during the printing and the frames can be retrieved as individual thermal images. The melt pool zone is defined as the region where the temperature exceeds the material’s melting point. For Inconel 718, this temperature threshold is $1336^{\circ}$ Celsius. The boundary of the melt pool surface is represented by an isotherm, as illustrated in Figure 3. Each pixel in the thermal images represents the corresponding temperature of the melt pool zone and the rest of the area within the field of view of the camera. Then, we perform a $360^{\circ}$ X-ray Computed Tomography (XCT) (SkyScanner1272) scan on Sample 2 (from Table 2) to identify the locations of the inter-layer pores and label the corresponding melt pool images captured at those specific locations. We find a total of 76 melt pools associated with internal pores and label them as class ‘1’ and the rest 1371 normal melt pools as class ‘0’. Examples of normal and abnormal melt pool thermal images are shown in Figure: 3.
@@ 3.3.2 MAE Encoder classifier
As we already mentioned earlier, the MAE Encoder has a similar architecture as a ViT encoder (Figure 5) except for the MLP classifier head. Thus, we add an extra two-layer MLP classifier head to the end of the MAE Encoder model to perform the classification task here. Another difference, unlike the ViT classifier, the MAE model lacks a designated class embedding ($[CLS]$). Thus, we apply max-pooling to the output embeddings and generate a similar class representation (Algorithm 3). We train this setup of fine-tuned MAE Encoder with the MLP classifier extension on the labeled melt pool data. This additional MLP head allows the model to map learned features to the desired class labels, and perform classification. The goal is to evaluate the performance of the fine-tuned MAE Encoder paired with an MLP classifier, leveraging the features learned during the self-supervised fine-tuning phase on unlabeled melt pool images, trained on labeled data to classify melt pool categories effectively. The training steps are demonstrated in the Algorithm 3.
@@ 4.1 Training setup of self-supervised MAE
First, we fine-tune the pre-trained MAE model using 7812 unlabeled melt pool images collected from Sample 1 from Table 2 in a self-supervised manner to learn the spatial features through image reconstruction. The data is divided into an 80%-20% split for the training and validation. We use the MAE-base model [32], configured with an encoder consisting of 12 Transformer blocks with a width of 768, and a decoder with 8 Transformer blocks and a width of 512 and reconstruction target without pixel normalization [27]. The original $400\times 400$ RGB images are resized to $224\times 224$ to reduce computational cost and match the dimensions of the pre-trained MAE. To make the model robust, data augmentation techniques such as rotation and flipping are applied, and drop-out and weight decay are used for regularization to avoid over-fitting. The training setup and specific hyperparameters are summarized in Table 3 and the learning curve obtained is demonstrated in Figure 6.
@@ 4.2 Training setup of supervised Classifiers
At this stage, we employ two different supervised classifiers, i.e., the fine-tuned MAE Encoder only with an MLP classifier head and a supervised ViT Classifier model. The reason behind attempting the MAE Encoder with MLP head is to test the ability of the MAE Encoder for the recognition task which is proven to be efficient on benchmark image datasets [27]. We also employ a fully supervised ViT Classifier base model [33]. We transfer the fine-tuned parameters of the self-supervised MAE Encoder to both classifiers. The same computational resources outlined in section 4.1 are used for the training. We run this setup on 1447 labeled melt images collected from Sample 2 from Table 2. The dataset is initially divided into a test set comprising 15% of the total data, and the remaining 85% is used for training. During the training phase, we employ a six-fold cross-validation approach to ensure robust training and performance evaluation of our models. Within the cross-validation, the data was split into six equal folds, where in each iteration, one fold served as the validation set while the remaining five folds were used for training. This process is repeated six times. We use the same set of hyperparameters in both classifiers’ setups and perform end-to-end supervised training on the labeled dataset to perform the binary classification of the normal and abnormal melt pools. The training setup and model specifications are mentioned in the Table 4.
@@
One of the major differences between these two aforementioned classifier models is unlike the ViT classifier, the MAE Encoder model does not have a classifier head by default, thus we add a two-layer MLP head to serve this purpose. Another difference is MAE Encoder does not have the class embedding and token $[CLS]$ added to the visible input patch embeddings; a key feature in the ViT classifiers. This $[CLS]$ token represents an aggregated representation of the image, as it has attended to all the other tokens of image patches through the multi-head self-attention mechanism [26]. Due to the absence of the class token in the MAE Encoder, we add a Max-pooling [34] to output tokens to create a single representation in Algorithm 3, that serves as the input to the MLP classifier similar to $[CLS]$ of standard ViT classifiers.
@@ 4.3 Performance Evaluation
The slightly lower performance of the MAE Encoder Classifier can be attributed to its default design, which is primarily specialized for spatial feature extraction by reconstruction rather than classification. Thus, as mentioned earlier, the MAE model does not require a class embedding or an MLP head. While we address this by applying max-pooling to the MAE Encoder’s outputs to approximate a global representation and add an MLP head, this approach does not fully replicate the $[CLS]$ embedding’s role in the ViT model. In ViT, the $[CLS]$ embedding is shared across all transformer layers, allowing it to integrate all patches’ information, and enhancing its capacity for classification tasks. Thus, while both classifiers are initialized with fine-tuned MAE parameters, the ViT Classifier achieves improved performance in supervised classification compared to the MAE Encoder Classifier approach.

### v2

@@ 1 Introduction
In recent years, significant efforts have been made to develop in-situ monitoring and characterization methods for melt pool dynamics in laser additive manufacturing (LAM), especially for identifying internal defects within directed energy deposition (DED). [7] introduced a real-time porosity detection framework using Self-Organizing Map (SOM) clustering on thermal melt pool images, achieving 96% accuracy. Similarly, [8] proposed a fusion approach combining PyroNet (CNN) and IRNet (LRCN) for porosity prediction via decision-level integration of pyrometer and infrared images. [9] reached 98.44% accuracy using thermal history and morphological features with a K-Nearest Neighbor (KNN) model. Meanwhile, [10] employed a variational autoencoder (VAE) with Gaussian mixture modeling (GMM) and K-means clustering to detect melt pool anomalies in DED, reporting 94.52% accuracy. In laser wire DED, [11] leveraged YOLO-based CNNs for real-time segmentation and analysis, improving monitoring precision. [12] used a high-dynamic range camera with a KNN classifier to classify stability zones, achieving a 13% error rate. In the mean time, multi-sensor fusion and data-driven models have also shown promise. [13] integrated multiple sensors with machine learning models like SVM, KNN, Random Forest, and MLP, achieving a 90% true positive rate while [14] applied an unsupervised online learning method in laser metal deposition (LMD), reaching 76% accuracy with K-means and 97% with SOM for classifying images as healthy” or anomalous.” Another work by [15] introduced a defect diagnosis system in DED using a degree of irregularity (DOI) feature index, achieving up to 96.92% accuracy for balling defects, outperforming STFT-based CNN models. More recent developments continue to push the boundaries of in-situ monitoring accuracy. [16] proposed FixConvNeXt, achieving 99.1% accuracy in real-time melt pool state identification for L-DED while reducing computational costs. [17] demonstrated an acoustic-based approach using CNNs for defect detection in DED, reaching 89% overall accuracy and 98% AUC-ROC for keyhole pore detection. [18] introduced a Gaussian SVM classifier for layer-wise anomaly detection in DED, achieving up to 96.38% accuracy and a 95.34% F-score, surpassing benchmarks in efficiency. [19] utilized DenseNet-39 with high-speed camera images, achieving 99.3% accuracy in classifying melt pool states as stable or unstable, reducing computational demands. [20] employed a ConvLSTM autoencoder with high-speed camera video data for detecting defects like wire dripping and arcing, effectively enabling real-time anomaly detection in DED. We summerize the different learning methodologies and models employed for defect detection tasks through in-situ melt pool data within the DED process in Table 1.
@@ 2 Experimentation & Data collection
For data collection, we conduct the experiments using a customized powder-based DED system (AMBIT Core DED, Hybrid Manufacturing Technology, TX, USA). We mount a thermal imager Pyrometer camera (ThermaViz® TV200, Stratonics) inside the printing chamber for in-situ monitoring and melt pool data capture during printing. Since our DED setup is hybrid, including both additive and CNC machines, we do an off-axis installation of the camera at a viewing angle of 57 (as shown in Figure 2) to maintain a clear line of sight to the melt pool while avoiding obstruction from the laser head and deposition tool and ensuring the required working distance of $245\pm 30$ mm. The focal length is fixed at 42 mm. We perform a positional and temperature calibration of the pyrometer camera before each experiment. A Neutral density filter is also calibrated and used to prevent pixel saturation during the in-situ recording by the pyrometer. In this study, we print two single-track thin-wall Inconel 718 samples, each measuring 40 mm in length and consisting of 10 layers in Figure 2, with the scanning direction reversed on every layer to create a linear reciprocating scan pattern. The variations in printing strategies and process parameters are summarized in Table 2 below. Note that, we have fixed the exposure time of the pyrometer at 0.3 ms optimized by experimentation for both sample printing. The camera records the distribution of thermal gradient of the traveling melt pool during the printing and the frames can be retrieved as individual thermal images. The melt pool zone is defined as the region where the temperature exceeds the material’s melting point. For Inconel 718, this temperature threshold is $1336^{\circ}$ Celsius. The boundary of the melt pool surface is represented by an isotherm, as illustrated in Figure 3. Each pixel in the thermal images represents the corresponding temperature of the melt pool zone and the rest of the area within the field of view of the camera. The transmissivity of Inconel 718 is considered negligible as the material is highly opaque. The two-color pyrometer reduces emissivity effects by measuring infrared radiation at two wavelengths, with pre-calibration using a manufacturer-provided lamp and slope correction adjustments as needed. Then, we perform a $360^{\circ}$ X-ray Computed Tomography (XCT) (SkyScanner1272) scan on Sample 2 (from Table 2) to identify the locations of the inter-layer pores and label the corresponding melt pool images captured at those specific locations. We find a total of 76 melt pools associated with internal pores and label them as class ‘1’ and the rest 1371 normal melt pools as class ‘0’. Examples of normal and abnormal melt pool thermal images are shown in Figure: 3.
@@ 3.3.2 MAE Encoder classifier
As we already mentioned earlier, the MAE Encoder has a similar architecture as a ViT encoder (Figure 5) except for the MLP classifier head. Thus, we add an extra two-layer MLP classifier head to the end of the MAE Encoder model to perform the classification task here. Additionally, while the ViT classifier incorporates a class embedding and a designated token $[CLS]$ within the input patch embeddings, the MAE Encoder instead utilizes a dummy token during pretraining, which serves a similar function but is primarily intended for alignment with the ViT architecture rather than explicit classification. As advised in the original MAE paper [27], applying global pooling over the encoder’s output tokens is recommended to obtain a single image representation for classification. In our implementation, we followed this recommendation and evaluated different pooling strategies. We found that max-pooling [32] of the output features yielded superior performance. Thus, we apply max-pooling to the output embeddings and generate a similar class representation (Algorithm 3).
We train this setup of fine-tuned MAE Encoder with the MLP classifier extension on the labeled melt pool data. This additional MLP head allows the model to map learned features to the desired class labels, and perform classification. The goal is to evaluate the performance of the fine-tuned MAE Encoder paired with an MLP classifier, leveraging the features learned during the self-supervised fine-tuning phase on unlabeled melt pool images, trained on labeled data to classify melt pool categories effectively. The training steps are demonstrated in the Algorithm 3.
@@ 4.1 Training setup of self-supervised MAE
First, we fine-tune the pre-trained MAE model using 7812 unlabeled melt pool images collected from Sample 1 from Table 2 in a self-supervised manner to learn the spatial features through image reconstruction. The data is divided into an 80%-20% split for the training and validation. We use the MAE-base model [33], configured with an encoder consisting of 12 transformer blocks with a width of 768, and a decoder with 8 Transformer blocks and a width of 512 and reconstruction target without pixel normalization [27]. The original $400\times 400$ RGB images are resized to $224\times 224$ to reduce computational cost and match the dimensions of the pre-trained MAE. To make the model robust, data augmentation techniques such as rotation and flipping are applied. Hyperparameters are selected based on the empirical tuning to ensure stable convergence. A learning rate of $0.0001$ with the Adam optimizer balanced convergence speed and stability, preventing sharp oscillations. Weight decay $(0.00005)$ reduces overfitting by penalizing large weight magnitudes, encouraging smoother representations, while dropout $(0.5)$ enhances regularization by randomly deactivating neurons, preventing reliance on specific features. The training setup and specific hyperparameters are summarized in Table 3 and the learning curve obtained is demonstrated in Figure 6.
@@ 4.2 Training setup of supervised Classifiers
At this stage, we employ two different supervised classifiers, i.e., the fine-tuned MAE Encoder only with an MLP classifier head and a supervised ViT Classifier model. The reason behind attempting the MAE Encoder with MLP head is to test the ability of the MAE Encoder for the recognition task which is proven to be efficient on benchmark image datasets [27]. We also employ a fully supervised ViT Classifier base model [34]. We transfer the fine-tuned parameters of the self-supervised MAE Encoder to both classifiers. The same computational resources outlined in section 4.1 are used for the training. We run this setup on 1447 labeled melt images collected from Sample 2 from Table 2. The dataset is initially divided into a test set comprising 15% of the total data, and the remaining 85% is used for training.
During the training phase, we employ a six-fold cross-validation approach to ensure robust training and performance evaluation of our models. Within the cross-validation, the data was split into six equal folds, where in each iteration, one fold served as the validation set while the remaining five folds were used for training. This process is repeated six times. We use the same set of hyperparameters in both classifiers’ setups and perform end-to-end supervised training on the labeled dataset to perform the binary classification of the normal and abnormal melt pools. The training setup and model specifications are mentioned in the Table 4.
@@ 4.3 Performance Evaluation
The slightly lower performance of the MAE Encoder Classifier can be attributed to its default design, which is primarily specialized for spatial feature extraction by reconstruction rather than classification. Since the Vision Transformer (ViT) was originally designed for classification tasks, fine-tuning naturally aligns with its pre-existing architecture, allowing for efficient adaptation. In contrast, the MAE encoder was pre-trained for a reconstruction-based objective, requiring a significant shift in representation learning to accommodate classification. This transition may have resulted in suboptimal adaptation, as the encoder had to reconfigure its learned feature representations for discriminative rather than generative tasks. Consequently, the ViT classifier fine-tuned more effectively, while the MAE encoder required additional adjustments to integrate the newly introduced classification head, potentially leading to slower convergence and slightly lower performance.